Type of improvement Example(s) from the RepRap case study Added functionality •The ability to remove material in addition to adding it. The ability to mix multiple materials. Use of ceramics and pastes instead of thermoplastics. •Embedding wire and conductive materials. •Ability to combine additive and subtractive manufacturing (Hydra-MMM) Improved existing functionality •Faster, more efficient, more detailed and/or stronger output. Increased ease of assembly and use •Many design improvements incorporated in the second version Mendel over the older Darwin design. •Derivative designs such as Makerbot and Ultimaker Protobox. Lower cost •Design of an alternative belt drive mechanism. •Allowing the use of cheap skate bearings More suitable components (e.g. easier-to-acquire) •A drive system based on ubiquitous ball-chains as an alternative to industrial timing belt and pulleys. •The Sanguino was developed as an alternative to the Arduino. Specialization towards a certain application •Digital pottery system •Plant growth modeling device Interoperability with other systems •Compatibility with G-Code common in industrial CNC installations. •Writing platform independent software. •Adding USB interfaces and having the machine work independently from removable storage media. Improved design architecture (e.g. modularization, part consolidation) •Adoption of industrial standards for NC-machine control. •Change from multiple independent microcontroller in a token ring to a single master microcontroller architecture with an optional slave extruder controller. Refining operating techniques •Sharing better settings for making parts easier to separate from their support structure (if applicable) Improved sharing infrastructure •Thingiverse.com was developed as website to facilitate sharing of digital designs for physical objects and currently hosts over 3000 user contributed objects which include, documentation, discussions and the data to manufacture them or make derivative designs. •Adoption of Wikis and blogs for knowledge sharing. •Local user groups